OpenStudy (anonymous):

The function f(x) = 4(4)x represents the growth of a fly population every year in a remote swamp. Jackie wants to manipulate the formula to an equivalent form that calculates three times a year, not just once a year. Which function is correct for Jackie's purpose, and what is the new growth rate? options: f(x) = 4(4)^x; growth rate 400% f(x) = 4(4)^3x, growth rate 4% f(x) = 4(1.59)^3x; growth rate 59% f(x) = 4(1.59)^x; growth rate 4%

OpenStudy (mathmale):

I think you mean f(x)=4(4)^x, not f(x)=4(4)x. Small but important difference!

OpenStudy (mathmale):

\[y=4(4)^{x}\]

OpenStudy (mathmale):

The " ^ " symbol signifies exponentiation.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

hello @SolomonZelman

OpenStudy (solomonzelman):

Hello

OpenStudy (anonymous):

where ya at @mathmale

OpenStudy (anonymous):

i need yo help @phi

OpenStudy (anonymous):

@phi @phi @phi @phi @phi @phi @phi @phi

OpenStudy (solomonzelman):

For example, the following problem: The function \(\large\color{#003366}{\displaystyle f(x)=2(10)^x }\) represents the growth of a fly population every year in a remote swamp. John wants to manipulate the formula to an equivalent form that calculates four times a year, not just once a year. Which function is correct for John's purpose, and what is the new growth rate? \(\large\color{#003366}{\displaystyle f(x)_{\rm ~[once ~a ~ year]}=2(10)^x }\) \(\large\color{#003366}{\displaystyle f(x)_{\rm ~[4~times ~a ~ year]}=2(\sqrt[4]{10})^{4x} }\) Now, each x is 4 times a year, and the 4th root is there to keep the function equivalent. Approximately, \(\large\color{#003366}{\displaystyle f(x)_{\rm ~[4~times ~a ~ year]}=2(1.78)^{4x} }\) Now, the percent rate! You are growing, times 1.78 every time. That is equivalent to saying that each time is 178% of the previous, or in other words, you are growing by 78%.

OpenStudy (solomonzelman):

I meant that eadh x is 1/4th of a year...

OpenStudy (mathmale):

Nice work, Solomon! Stress that this was an EXAMPLE that your student needs to apply to his or her original problem statement.

OpenStudy (solomonzelman):

Thank you!